### v1

# Active nematic pumps

Abstract
Microfluidics involves the manipulation of flows at the microscale, typically requiring external power sources to generate pressure gradients. Alternatively, harnessing flows from active fluids, which are usually chaotic, has been proposed as a paradigm for the development of micro-machines. Here, by combining experimental realizations and simulations, we demonstrate that the addition of triangular-shaped inclusions into an active nematic gel can locally break the fore-aft symmetry of active turbulence and stabilize flow fields with self-pumping capabilities. The proposed strategy has enabled us to generate wall-free and self-powered microfluidic systems capable of both cargo transport and mixing along with the downstream flow. We analyze the performance of these active pumps, both isolated and within cooperative ensembles in terms of their output velocity and hydrostatic pressure buildup. Finally, we demonstrate strategies to incorporate them into specifically designed microfluidic platforms to advantageously tailor the geometry of active flows. Our results reveal new possibilities for leveraging the self-organized mechanodynamics of active fluids.
I Introduction
Microfluidic systems have become fundamental tools in modern science and technology, enabling precise manipulation of fluids at the microscale for applications ranging from chemical synthesis and microfabrication to biomedical devices [1]. While these systems have been extremely useful in a wide range of scientific and industrial fields, both their geometrical constraints and the fact that they depend on external pressure sources impose limitations on systems design, downscaling, and autonomy.
A transformative approach to microfluidics could be explored by transitioning from traditional to active microfluidic systems, leveraging the flows generated by self-powered and self-organized biological or bioinspired active fluids [2, 3, 4], as well as their intrinsic control mechanisms, refined during millions of years of natural evolution. Constituted by both synthetic particles [5] and self-propelled entities of biological origin, such as bacteria [6, 7, 8], mamallian cells [9, 10], or cytoskeletal proteins [11, 12], active fluids efficiently convert chemical energy into sustained motion, typically in the form of seemingly disordered flows [13, 14, 15, 16]. Understanding how to regulate the flows of these active fluids [17, 18] can thus be a handle to harnessing their mechanodynamics into well-organized flow patterns for the development of new microfluidic technologies.
Within the state-of-the-art active systems, synthetic active gels stand out as exceptional materials [19]. They combine the best attributes of biology and colloidal sciences, offering not only efficient self-organized behaviors and adaptability but also robustness and ease-of-manipulate. In particular, active nematic (AN) gels, endowed with orientational order, provide additional unique properties that enhance their functionality and controlability [20]. In this regard, some promising steps toward controlling ANs have been realized in the past few years. For instance, it has been experimentally demonstrated that spatial confinement within channels or rings can lead to an effective rectification of turbulent flows in active nematic preparations, both in 3d and in 2d, resulting in regimes with effective transport [17, 21, 22, 23]. Flow control of active fluids can also be exerted using interfacial conditioning [24, 25], or light patterning [26]. The symmetry breaking that leads from isotropic active turbulence to net active transport needs a mechanism to locally set both the orientation and the handedness of the flows. This has been typically addressed by using ratchet-like boundaries [27, 17, 23], which have enabled to rectify the otherwise random rotation of inclusions embedded in an active fluid [2, 28, 4]. In spite of these developments, we still lack a versatile strategy to unleash the full microfluidic potential of active nematic flows.
Figure 1: Active pumping with a square array of triangular inclusions. a An array of hydrogel columns with triangular cross section is photopolymerized in-situ in the aqueous phase of an AN flow cell. Scale bar 100 $\mu$m. b Fluorescence image showing the adaptation of the AN to the presence of the obstacle array (see Movie Active nematic pumps). c Simulated system (see Movie Active nematic pumps). The instantaneous director field is rendered using the Line Integral Convolution filter. Diagrams d and f correspond to the analysis of the experimental flow field, while e and g correspond to the analysis of the simulated flow field. Experimental images correspond to the central part of a lattice containing 6 rows and 9 columns in total, while the simulated system, also only partially shown, contains 6 rows and 30 columns with periodic boundary conditions in both directions. Time-averaged velocity (d, e) and vorticity (f, g) fields. $v_{r}$ is the average speed and, in the simulations, $\omega_{r}=v_{r}/D_{def}$, where $D_{def}$ is the average distance between defects, all measured in the absence of obstacles. h Simulated average horizontal flow velocity relative to $v_{r}$ as a function of the size of the equilateral triangle, $H$, for three different normalized activity coefficients, $\zeta D_{def}/(v_{r}\eta)$, and for a lattice periodicity $L=128$ l.u. (or $\rm 2.3\,D_{def}$). Here, $\eta$ is the two-dimensional active fluid viscosity. Simulations are performed in one lattice cell, using periodic boundary conditions. Data are collapsed by rescaling $H$ with $D_{def}$. i Diagram of the normalized average horizontal velocity, $v_{x}/v_{r}$, as a function of the normalized triangle size, $H/D_{def}$, and triangle size relative to cell size, $H/L$. Simulations are performed with $\zeta D_{def}/(v_{r}\eta)=7.84$.
Here, we address this gap by demonstrating that inclusions with well-defined geometrical asymmetry can lead to the phenomenon of active pumping. In particular, we take advantage of a recent protocol to polymerize biocompatible hydrogel structures directly into a fully-developed active nematic gel [29] to imprint lattices of triangular inclusions that, for sizes comparable to the intrinsic active length scale, lead to hydrostatic pressure buildup and enhanced unidirectional flows. Combining experiments and numerical simulations we show that the observed phenomenon is a cooperative effect emerging from a lattice of counter-rotating vortices that result from the presence of obstacles with broken fore-aft symmetry within the active nematic fluid. By characterizing the interplay between flow patterns and pressure gradients generated within the active nematic fluid, we demonstrate that arrays of triangular inclusions can lead to efficient active nematic pumping through the creation of virtual flow channels. The absence of physical confinement walls makes this system an ideal candidate for large-scale active microfluidic transport and mixing of colloidal cargo, or to locally adjust the geometry of the active fluid [30].
Figure 2: Active pumps in series and in parallel. a Simulated study of the performance of pumps in series. The number of pumps acting in parallel (number of flow channels) is kept constant at 3, while the number of columns, $N$ is varied. The yellow square represents a single pump in the region between four triangles. The size of the simulated domain, $36\times 6$ cells (or $\rm 82.3\times 13.7\,D_{def}$), and the simulation conditions, $H=1.25D_{def}$, $L=2.3D_{def}$, $\zeta D_{def}/(v_{r}\eta)=7.84$ and closed box with no-slip boundary conditions, are the same in all cases. b Velocity field along the central lines between rows of triangles (averaged for the three rows and with time) for different numbers of pumps in series, including $N=0$ (closed box without triangles) for comparison. The orange dashed vertical line indicates the exit of the pump region. c For the same simulations as in (b), we monitor the pressure field. Here we plot the normalized pressure profile, averaged for the three rows and with time, and normalized with $p_{r}=\rho_{0}v_{r}^{2}$, where $\rho_{0}$ is the average density. d Average normalized horizontal velocity at the output of the pump system for different numbers of pumps in series. The inset indicates the corresponding pressure difference across the pump system. The error bars represent the standard deviation for different random initial conditions. e System of pumps in parallel. The number of pumps acting in series is kept constant at 9 (10 columns of triangles), while the number of rows, $N$, is varied. f Average flow rate at the output for different numbers of pumps in parallel, relative to the reference flow rate $Q_{r}=v_{r}D_{def}$. The inset depicts the pressure difference across the pump system.
II Results
II.1 Creation of active nematic pumps
We photopolymerize square lattices of hydrogel columns with triangular cross section inside an AN flow cell [29]. The columns are attached to the hydrophilic glass plate and extend across the AN towards the thin oil layer on the complementary hydrophobic plate (see Fig. 1a and Methods). As a result, a lattice of rigid triangular obstacles is suddenly embedded within the AN, thus forcing the adaptation of its orientational and velocity fields to the new in-plane boundary conditions (Fig. 1b).
Both experiments and continuum hydrodynamic simulations (Fig. 1c; see Methods) show that the active flow patterns organize around the array of obstacles leading to the formation of virtual flow channels. The velocity field is oriented, on average, along the space between rows of obstacles, antiparallel to the orientation of the triangles in the array (left-to-right in the images), as shown in Fig. 1d and Fig. 1e. This is different from active nematic flows confined in channels, which can exhibit net transport of the normally turbulent-like flows, but without a predefined flow polarity [17, 22]. In particular, we observe that active flows organize as a pair of counter rotating vortices on both sides of the frontal tip of each triangle. This is clearly highlighted in Fig. 1f and g, where we represent the average vorticity field of the active flows. Our observations are consistent with earlier simulations that reported a steady vortex lattice forming in AN confined in an array of astroid-shaped obstacles [31]. In our case, the flows emerging from counter rotating vortex pairs leads to net transport in a direction that is predefined by the orientation of the obstacles in the array. Inside the virtual channels, flow speed is non monotonous, being higher between vortex pairs. Overall, this leads to a unidirectional pumping mechanism. (Fig. Active nematic pumps). A similar flow rectification effect has been recently proposed using numerical simulations with half-astroid-shaped obstacles [32], consistent with the expected ratchet effect in the presence of asymmetric obstacles.
The reported pumping mechanism depends on the number, size, shape, and spatial distribution of the triangles in the array. A stable vortex lattice emerges only if at least 4 inclusions are present (Fig. Active nematic pumps). We consider a 2 by 2 square lattice of triangular inclusions to be the minimal system to achieve active pumping. In Fig. 1, the size and spacing of the triangles are chosen to be similar to the intrinsic length scale of the AN, $D_{def}$, which we have determined to be of the order of 70 $\mu$m from the average spacing between defects in the absence of obstacles (see early stages of Movie S1). Using numerical simulations, we have studied the effect of these two parameters.
In Fig. 1h, we show the average normalized horizontal velocity as a function of the normalized size of the triangles, $H/D_{def}$, for a fixed spacing $L=2.3D_{def}$. Note that, in the absence of AN pumping effects, the flows are turbulent-like and therefore, the net velocity is zero in any direction. We observe a quick increase of the pumping efficiency with obstacle size, until $H\simeq D_{def}$, slowly decreasing for larger obstacles. We have performed simulations for different activity coefficients (related to the concentration of adenosine triphosphate in experiments), and find that the results remain invariant as the velocity and spatial dimensions are rescaled with the intrinsic velocity and active length scale in the absence of obstacles at each activity [33, 20].
We next use simulations to explore, for a constant activity, the optimal geometry of the square lattice of triangles, in terms of their size and spacing. The resulting phase diagram, shown in Fig. 1i, indicates that the optimal arrangement corresponds to $H\simeq D_{def}$ and $L\simeq 2D_{def}$, which we have used in the experiments and simulations in Fig. 1, and will be used henceforth by default. Our experimental technique allows to fabricate more compact active pump arrangements, as shown in Fig. Active nematic pumps, thus demonstrating the capabilities for miniaturization of these devices.
The pumping effect is also severely affected by the lattice regularity, as analyzed with simulations in Fig. Active nematic pumps for lattices of triangles with different degrees of disorder with respect to the perfect square lattice arrangement. We observe that a random lateral displacement of the triangles’ center 25% with respect the square lattice nodes halves the efficiency of the AN pumping effects, which are completely lost when lateral displacement is higher than 50 %.
II.2 Active pump systems
After determining the optimal geometric features of basic AN pumps, we consider the design of more intrincate microfluidic systems. In particular, we have studied the performance of square lattices of triangles (Fig. 2) for different number of columns (active pumps in series, Fig. 2a-d) and rows (active pumps in parallel, Fig. 2e,f). We observed that, by concatenating AN pumping units, the average speed increases with the number of columns (Fig. 2b), until reaching a maximum value at about 10 pumps in series (Fig. 2d).
In parallel, we also consider the concurrent profiles of the hydrostatic pressure within the active fluid along the flow channels, and we observe that it increases steadily with the number of pumping elements in series, even when the average speed has reached its saturation value (Fig. 2c and inset in Fig. 2d). In our active pump system, pressure gradients are positively correlated with the flow velocity. This is contrary to flow through a microfluidic channel or through a porous medium, where velocity is negatively correlated with the driving pressure gradients, even in the case of pressurized active fluids [34]. In classical hydraulics, a system of ideal pumps in series yields a flow velocity that does not change with the number of pumps and a pressure that builds up linearly with the number of elements [35]. Interestingly, our active pumps behave as ideal hydraulic pumps when the number of elements in series is larger than 10, as observed in Fig. 2d.
[…]
Figure 4: Microfluidic applications of active nematic pumping. a Experiments where the output of an active pump system is injected into a channel. Triangular obstacles and channel boundaries are overlaid on the map of the velocity field. Scale bar 100 $\mu$m. b Horizontal velocity (averaged over the central $50\,\mu$m along the channel) relative to $v_{r}$. c Simulation of an active pump system injecting into a channel of width $L_{c}=2.3D_{def}$ and activity $\zeta D_{def}/(v_{r}\eta)=7.84$. We perform simulations inside a closed box of dimensions $24\times 6$ cells (or $\rm 54.9\times 13.7\,D_{def}$), with no-slip boundary conditions on the walls. d Velocity field along the central part of the channel, averaged in time and for the different initial conditions in simulations, for three different normalized channel widths $L_{c}/D_{def}$. e Velocity field inside the channel with and without the pump system for different channel widths in simulations. The error bars are obtained from the different initial conditions. f Simulations of larger scale flows guided by a lattice of triangles. Clockwise flow circulation is guided by the active pump system. The white and the grey trajectories illustrate the path followed by two tracers and the white rectangle indicates the path used to calculate the circulation. The system, with dimensions $16\times 10$ cells (or $\rm 36.6\times 18.3\,D_{def}$), is in a closed box with solid walls at the borders. g Circulation, $C=\oint\mathbf{u}\cdot\mathbf{dl}$, vs. time along the rectangular circuit in the simulations in f. h Experimental time-averaged velocity field in the presence of a closed circuit of elementary active pumps. Scale bar $100\mu$m.
Next, we analyze the performance of active pumps in parallel, as shown in Fig. 2e. In this case, the setup includes a fixed number of 10 columns of triangles (9 pumps in series), and we study the performance of the system when changing the number of rows. The output velocity is the one expected from 9 pumps in series, so we will characterize each pump system with its average flow rate, which we find to be proportional to the number of elements in parallel, just as it would be for a system of ideal hydraulic pumps (Fig. 2f). We have also monitored the average pressure drop across the system, which we find to be constant, independent of the number of elements in parallel. These observations show that neighboring pumping elements along the same column operate independently, as a system of ideal pumps, despite the fact that the separation between adjacent flow channels is permeable. This favors not only net transport but also matter exchange, as we show in the next section.
II.3 Micro-fluidic transport and mixing by active nematic pumps
Next we demonstrate how to harness active flows to achieve effective microfluidic transport and mixing. This is implemented by locally imprinting our self-pumping system that, being realized along virtual channels, readily enables exchange of matter across their discontinuous walls. Both experiments (Fig. 3a,b) and simulations (Fig. 3c,d) reveal that passive colloidal particles are advected downstream by active flows, despite following erratic trajectories. Indeed, particles can be trapped in a vortex or they can switch to neighboring channels in the parallel pump system, but with a net average displacement clearly oriented downstream (Fig. 3b,d). A detailed statistical study of particle velocities (Fig. Active nematic pumps) show that they reach a steady average value along the channels with the orientation set by the triangular obstacles. Conversely, the transversal ($V_{y}$) component has a vanishing average value, consistent with no net transversal flow, but the average value of $|V_{y}|$ is finite, indicating an active exchange of particles across the virtual channel walls.
On the other hand, channel permeability represents an enormous advantage with respect to classical microfluidic devices, and allowed us to achieve enhanced downstream mixing. We tested this concept by simulating an active pump system where adjacent circular ensembles of two distinctive colloidal species (red and blue in Fig. 3f) are advected by active flows. We observe that, while the center of mass of the ensemble moves downstream, mixing of the transported particles takes place, both in the longitudinal direction within channels and between neighboring channels. In fact, most particles move to a different channel during the process, as demonstrated in Fig. Active nematic pumps, where we show that the probability of a particle remaining in its original channel drops steadily with time. The effectiveness of mixing is exemplified in Fig. 3g with the temporal evolution of a mixing parameter, $\chi$, which drops from $\chi=1$ (unmixed) to $\chi\simeq 0.1$ (near perfect mixing) after the ensemble has crossed $\sim 11$ pumping elements. Although mixing is more efficient in fully turbulent active flows performing free of lateral constraints, our pumping system combines mixing with downstream transport, something that cannot be achieved through a lattice of isotropic obstacles or inside a channel with smooth walls (Fig. Active nematic pumps).
II.4 Integrating active pumps into microfluidic systems
The versatility of our active nematic pumping system enables its incorporation anywhere and at anytime within an active fluid layer. As an example, in Fig. 4 we demonstrate the efficacy of our control strategy by directing the flow from the outlet of an AN-pumped virtual channel into a channel confined by physical walls. It is known that flow inside a channel narrower than the active length scale can lead to structured flow patterns and even to net transport, although the resulting flow orientation is randomly selected upon symmetry breaking [23]. In the present case, the orientation of the flow is determined by the active pumping system. Here, both triangular inclusions for pumping and the continuous walls of the channel were fabricated simultaneously. One could, however, incorporate an active pump to feed a preexisting channel. In this system, we observe that the rectified flow speed decays downstream inside the channel, but it maintains the directionality, which is imposed by the active pumps. In this example, the length of the channel is similar to that of the active pump system, and is limited by the field of view in our current experimental setup. We have confirmed the experimental results using numerical simulations with a channel much longer than the size of the AN pumping unit (Fig. 4c). We find that the average velocity in the center of the channel stabilizes to the value at the output from the pump (Fig. 4c). In these simulations, the channel is just slightly wider than the flow lane of the injecting pump system. For wider channels, we observe a decay of the speed, similar to the one observed in experiments (Fig. 4d). Results are different if the output of the active pump system is directed towards an arrangement of non-pumping obstacles, such as an array of circular columns (Fig. Active nematic pumps). The permeable channel walls lead to a quick randomization of the velocity.
Finally, we designed a system that sustained a steady active flow circulation along channels of arbitrary shape. As an example, in Fig. 4f we have simulated a large rectangular loop of active pumps. The average clockwise flow circulation along the center of the channel (rectangular path in the figure) is shown to fluctuate around a steady value for the duration of the simulations (Fig. 4g). Interestingly, the active flows in the region surrounded by the loop of active pumps are no longer isotropic, which provides us with a tool to generate alternative flow patterns. For instance, in Fig. 4h, we show the average flow field created experimentally with a simple ring of elementary active pumps, revealing that the central region of the AN layer has developed a steady counter clockwise flow.
III Discussion
Achieving flow control is crucial in lab-on-a-chip devices. Fluids are driven by pressure gradients that propagate through miniaturized channels and valves with flows that obey the fundamental laws of hydraulics, namely, mass conservation and fluxes that are proportional to local pressure gradients. The strategies developed for these microfluidic devices cannot be directly adapted to harness the power of active fluids, where flows are self-sustained, often inherently disordered, and evolving following modified laws of hydraulics, as it has been recently demonstrated [36, 34].
Amidst this complexity, we have demonstrated here a versatile protocol to locally control the directionality of the flows of a microtubule/kinesin AN. By incorporating photopolymerizable reactants in the aqueous medium, we have polymerized lattices of columns with triangular cross section, which force an adaptation of the surrounding active flows, breaking the fore-aft symmetry and transforming the usual isotropic active turbulence regime into ordered arrays of counter-rotating vortices that organize the AN flows along virtual channels, as material flux is allowed across them. Combining experiments and numerical simulations, we have analyzed both the flow velocity and hydrostatic pressure profiles, demonstrating that the new structures behave as classical hydraulic pumps when a small number of elements are created sequentially, with a velocity and hydrostatic pressure that are proportional to the number of pumping elements.
The demonstrated active pump systems can be implemented to tailor complex active flow patterns, allowing to select both the direction and magnitude of the self-sustained flows, and can also be used to design arbitrary mixing patterns of advected particles. Overall, the proposed strategy harnesses the self-sustained flows generated by active nematics into active fluid pumping, possessing the necessary qualities for autonomous transport and mixing in future bioenabled microfluidic devices.
We have also demonstrated that a suitable design strategy will allow to force active fluids into developing flow patterns that would not normally occur as the result of their self-organization, such as large scale steady flow circulation. In summary, active nematic pumps can bring us closer to effectively harnessing self-sustained active fluid layers for the development of bio-inspired micromachines.
Finally, it is also worth stressing that the use of biocompatible fabrication protocols, as proposed here, ensures the general applicability of these results to different active fluids, both synthetic and of biological origin.
Acknowledgements
We thank M. Pons, A. LeRoux, and G. Iruela for their assistance in the expression of motor proteins. I.V.-C. acknowledges funding from Generalitat de Catalunya though a FI-2020 PhD. Fellowship. P.G. acknowledges support from the European Union’s Horizon Europe Research and Innovation Programme through the Marie Sklodowska-Curie Actions (grant agreement 101065794). I.V.-C., J.I.-M., and F.S. acknowledge funding from MICIU/AEI/10.13039/501100011033 (Grant No. PID2022-137713NB-C21). R. C. and M.T.G. acknowledge financial support from the Portuguese Foundation for Science and Technology (FCT) under the contracts: PTDC/FISMAC/5689/2020 (DOI 10.54499/PTDC/FIS-MAC/5689/2020), EXPL/FIS-MAC/0406/2021 (DOI 10.54499/EXPL/FIS-MAC/0406/2021), UIDB/00618/2020 (DOI 10.54499/UIDB/00618/2020), UIDP/00618/2020 (DOI 10.54499/UIDP/00618/2020), DL 57/2016/CP1479/CT0057 (DOI 10.54499/DL57/2016/CP1479/CT0057) and 2023.10412.CPCA.A2. The authors are indebted to the Brandeis University MRSEC Biosynthesis facility, supported by NSF MRSEC 2011846, for providing the tubulin.
Methods
III.1 Experiments
III.1.1 Active gel preparation
Microtubules (MTs) were polymerized from heterodimeric ($\alpha$,$\beta$)-tubulin from bovine brain (Brandeis University Biological Materials Facility). Tubulin (8 mg mL-1) was incubated at 37 °C for 30 min in aqueous M2B buffer (80 mM Pipes (piperazine-N,N’-bis(2-ethanesulfonic acid)), 1 mM EGTA (ethylene glycol-bis($\beta$-aminoethyl ether)-N,N,N’,N’-tetraacetic acid), 2 mM MgCl2) (Sigma; P1851, E3889 and M4880, respectively) prepared with Milli-Q water and supplemented with the antioxidant agent dithiothrethiol (DTT) (Sigma 43815) and GMPCPP (guanosine-5’-[($\alpha$,$\beta$)-methyleno]triphosphate) (Jena Biosciences, NU-405) up to a concentration of 1 mM and 0.6 mM, respectively. GMPCPP is a non-hydrolysable guanosine triphosphate (GTP) analogue that favours the formation of stable MTs, obtaining high-density suspensions of short MTs (1-2 $\mu$m). 3% of the total tubulin concentration was labelled with Alexa 647 to allow their observation by means of fluorescence microscopy. Afterward, the mixture was annealed at room temperature for 4 h and kept at -80 °C until use.
Drosophila melanogaster heavy-chain kinesin-1 truncated at residue 401 fused to biotin carboxyl carrier protein (BCCP), and labeled with six histidine tags (K401-BCCP-6His) was expressed in BL21(DE3)pLysS competent cells using the plasmid WC2 from the Gelles Laboratory (Brandeis University) and purified with a nickel column.After dialysis against 500 mM imidazole aqueous buffer, kinesin concentration was estimated by means of absorption spectroscopy. Finally, the kinesin was stored in a 40% (wt/vol) aqueous sucrose solution at -80 °C until use.
Kinesin motor clusters were prepared by incubating on ice for 30 min the motor protins with streptavidin (Invitrogen; 43-4301) at a specific ratio 2:1. MTs were mixed with the motor clusters, ATP (adenosine 5’-triphosphate) (Sigma; A2383) to fuel the motors, MTs, and the depleting agent PEG (poly(ethylene glycol)) (20 kDa, Sigma; 95172). An enzymatic ATP regenerator system was added to ensure the duration of activity for several hours, it consisted on phosphoenolpyruvate (PEP) (Sigma; P7127) that fuelled the enzyme pyruvate kinase/lactate dehydrogenase (PKLDH) (Sigma; P0294) to convert ADP (adenosine 5’-diphosphate) back to ATP. The mixture also contained several antioxidants to avoid photobleaching during the fluorescence imaging: DTT, trolox (6-hydroxy-2,5,7,8-tetramethylchroman-2-carboxylic acid) (Sigma; 238813), glucose oxidase (Sigma; GT141), catalase (Sigma; C40), and glucose (Sigma; G7021). Furthermore, the constituents of the photopolymerizing hydrogel were added to the active gel. This photopolymerizing hydrogel was composed of a photoinitiator, LAP (lithium phenyl-2,4,6-trimethylbenzoylphosphinate) (TCI; L0290), and a monomer, a multiarmed PEG derivative with acrylate groups at each terminal (4-ArmPEG-Acrylate 5 kDa) (Biochempeg; A44009-5k). Additionally, Dragon Green (DG) fluorescent polystyrene particles of diameter 1 $\mu$m (Bang Laboratories; FSDG004) were added to extract the flow field of the active material. Table 1 shows the final concentrations of each reagent used in the preparation of the active material.
Compound Buffer Final Conc. Units
Microtubules Original 1.33 mg/mL
Kinesin Original 0.32 $\mu$M
Streptavidin M2B 0.01 mg/mL
ATP M2B 300 $\mu$M
PEG (20 kDa) M2B 1.6 % w/v
PEP M2B 27 mM
MgCl2 M2B 3.3 mM
DTT M2B 5.8 mM
Trolox Phosphate 2.1 mM
Catalase Phosphate 0.04 mg/mL
Glucose Phosphate 3.20 mg/mL
Glucose Oxidase Phosphate 0.21 mg/mL
PK Original 27 u/mL
LAP Mixture 0.226 % w/v
4PEG5k Mixture 4.52 % w/v
DG Particles Original 1 % v/v
Table 1: Composition of all stock solutions (including buffer used for their preparation), and concentration of the different species in the final mixture. Acronyms used in this table are: PEG (Poly-ethylene glycol); PEP (Phosphoenol pyruvate); ATP (Adenosin triphosphate); PK (Pyruvate Kinase); DTT (1,4-dithiothreitol); 4PEG5k (4-ArmPEG-Acrylate 5 kDa); LAP (lithium phenyl-2,4,6-trimethylbenzoylphosphinate). The two latter compounds where directly incorporated to the active mixture. M2B buffer: 80 mM PIPES (piperazine-N,N’-bis(2-ethanesulfonic acid)) pH 6.8, 2 mM MgCl2 1 mM EGTA (egtazic acid). Phosphate buffer: 20 mM Phosphate buffer (6.68 mM KH2PO4, 12.32 mM K2HPO4) pH 7.2; Original: species is obtained already dissolved in its custom buffer.
III.1.2 Active nematic cell
Flow cells were prepared with a bioinert and superhydrophilic polyacrylamide-coated glass and a hydrophobic aquapel-coated glass. A 50 µm height double-sided tape was used as a spacer, leaving a channel of $\sim$4 mm width. The cell was filled first with fluorinated oil (HFE7500; Fluorochem; 051243) which contains a 2% of a fluorosurfactant (008 Fluorosurfactant; RanBiotechnologies) to get a surfactant-decorated interface. Afterwards, the active material was introduced into the cell by capillarity and the cell was sealed using petroleum jelly to avoid evaporation.
III.1.3 Imaging and photopatterning
The observation of the active material was done by means of fluorescence microscopy. We used a Nikon Eclipse Ti2-e equipped with a white LED source (Thorlabs MWWHLP1), a Cy5 filter cube (Edmund Optics) and a FITR filter cube (Thorlabs, TLV-TE2000-FITC). Images were captured using an Andor Zyla 4.2 Plus camera or a QImaging ExiBlue cooled CCD camera operated with ImageJ $\mu$-Manager open-source software.
The fluorescence microscope was modified in order to incorporate a Ti Light Crafter 4500 DLP development modules (EKB Technologies, Ltd.) equipped with a high 2 W 385 nm LED source in one case. Projected patterns are incorporated into the light path of the inverted microscope by means of a collimating lens (f = +150 mm) and a 505 nm dichroic mirror (Thorlabs DMLP505R), and are focused on the sample by means of the microscope objectives, reaching a lateral resolution up to a few microns. The DMD projector is connected as an external monitor to a computer, thus enabling real time control of the projected patterns using MS-PowerPoint slides. Hydrogel objects were polymerized using the x10 objective, resulting in a light power density of 1.6 W cm-2), and a irradiation time of 2 seconds.
III.1.4 PIV analysis
The flow field of the active material was extracted by means of particle image velocimetry (PIV) from the fluorescent images of the Dragon Green particles dispersed in the material using the public MatLab App PIVLab.
III.2 Simulations
We model the active nematics using a 2D continuum hydrodynamic theory which is summarized in this section. The director field $\mathbf{n}(\mathbf{x})$ gives the average orientation of the filaments in the AN at the position $\mathbf{x}$ while the degree of order is given by the scalar order parameter $S$, which is $S=0$ in the isotropic phase and has a finite value $S_{N}$ in the nematic phase. Because the filaments have a head-tail symmetry, a tensor order parameter is used: $Q_{\alpha\beta}=2S\left(n_{\alpha}n_{\beta}-\delta_{\alpha\beta}/2\right)$. In the free energy density, the bulk terms are not considered. It is assumed that the fluid becomes nematic only when it is active, which mimics the behaviour of the mixtures of microtubule and kinesin [37]. The free energy density includes an elastic term responsible for penalizing distortions in the order parameter, $f_{E}=K\left(\partial_{\gamma}Q_{\alpha\beta}\right)^{2}/2$, where $K$ is the elastic constant, and an anchoring term applied only at the solid surfaces, $f_{W}=W(Q_{\alpha\beta}-Q_{\alpha\beta}^{0})^{2}/2$, which penalises deviations from the surface-preferred order parameter $Q_{\alpha\beta}^{0}$ with strength $W$. The total free energy is the sum of these two terms integrated in the entire area: $\mathcal{F}=\int fd^{2}x$, where $f=f_{E}+f_{W}$ at the solid surface and $f=f_{E}$ elsewhere. The dynamics is governed by the Beris-Edwards (for the tensor order parameter $Q_{\alpha\beta}$), continuity and Navier-Stokes equations (for the velocity field $\mathbf{u}$), respectively [38, 39, 20]:
$\partial_{t}Q_{\alpha\beta}+u_{\gamma}\partial_{\gamma}Q_{\alpha\beta}-S_{\alpha\beta}=\Gamma H_{\alpha\beta},$ (1)
$\partial_{\alpha}u_{\alpha}=0.$ (2)
$\rho\partial_{t}u_{\alpha}+\rho u_{\beta}\partial_{\beta}u_{\alpha}=-\partial_{\alpha}p+2\eta\partial_{\beta}D_{\alpha\beta}-\zeta\partial_{\beta}Q_{\alpha\beta}.$ (3)
In Eq. (1), $\Gamma$ stands for the rotational diffusivity, $H_{\alpha\beta}$ for the molecular field: $H_{\alpha\beta}=-\frac{\delta\mathcal{F}}{\delta Q_{\alpha\beta}}+\frac{\delta_{\alpha\beta}}{2}\operatorname{Tr}\left(\frac{\delta\mathcal{F}}{\delta Q_{\gamma\epsilon}}\right)$ and $S_{\alpha\beta}$ for the co-rotational term: $S_{\alpha\beta}=(\xi D_{\alpha\gamma}+W_{\alpha\gamma})\left(Q_{\beta\gamma}+\delta_{\beta\gamma}/2\right)+\left(Q_{\alpha\gamma}+\delta_{\alpha\gamma}/2\right)(\xi D_{\gamma\beta}-W_{\gamma\beta})-2\xi\left(Q_{\alpha\beta}+\delta_{\alpha\beta}/2\right)(Q_{\gamma\epsilon}\partial_{\gamma}u_{\epsilon})$, where the vorticity and shear rate are $W_{\alpha\beta}=(\partial_{\beta}u_{\alpha}-\partial_{\alpha}u_{\beta})/2$ and $D_{\alpha\beta}=(\partial_{\beta}u_{\alpha}+\partial_{\alpha}u_{\beta})/2$. The aligning parameter $\xi$ controls the relative importance of the vorticity and shear rate and it is related to the shape of the particles that compose the nematics: rod-like particles have $\xi>0$, while disk-like particles have $\xi<0$. We consider rod-like particles (filaments) in the flow-aligning regime. In Eq. (3), $\rho$ is the fluid density, $p$ is the fluid pressure and $\eta$ is the absolute viscosity. The activity level is controlled by the parameter $\zeta$, which is positive for extensile systems as those in this work. The equations described above are a simplified version of a standard 2D model for active nematics [37, 40, 20]. Because the active stresses are typically much larger than the passive ones [29], we neglected the passive stress in Eq. (3) as in previous works [22, 41].
A hybrid numerical method is used in the simulations. Equation (1) is solved using a predictor-corrector finite-differences method while Eqs. (2) and (3) are recovered in the macroscopic limit using the lattice Boltzmann method [42, 40]. At the solid boundaries, we use planar weak anchoring and no-slip boundary conditions by means of half-way bounce-back boundary conditions. The initial conditions are the fluid at rest and directors mostly aligned with the horizontal direction with random fluctuations of ±$\pm 2^{\circ}$. Different seeds are used to generate random initial conditions. The parameters used in the simulations are as follows (except where otherwise stated) in lattice units, in which the reference density, lattice spacing and time step are equal to one: absolute viscosity $\eta=6.67$, elastic constant $K=0.015$, aligning parameter $\xi=0.9$, activity $\zeta=0.01$, fluid density $\rho=40$, anchoring strength $W=0.002$, rotational diffusivity $\Gamma=0.4$, obstacle size $H=70$ and size of the square unit cell $L=128$.
The results are shown in reduced units, which are calculated considering three quantities in lattice units: the average density $\rho_{0}=40$, the average velocity of the free active nematics (without obstacles) $v_{r}\approx 0.0105$ and the average distance between defects for the free active nematics $D_{def}\approx 56$. All other reference quantities are calculated using these three.